# Constraint-Aware Zero-Shot Vision-Language Navigation in Continuous Environments

### v1

@@ II Related Work
Although significant progress has been made, these methods overlook low-level control in continuous environments, particularly in scenarios where predefined navigation graphs are unavailable. As a result, their performance drops significantly when VLN methods are applied to real-world settings. [43].
@@ III-B Constraint-aware Sub-instruction Manager
Instruction Decomposition. As shown in Figure 2, at the beginning of an episode, CSM decomposes the instruction into a sequence of sub-instructions. Each of them outlines the goal for the current sub-instruction and specifies the constraints for switching to the next one. We categorize these constraints as object constraints (e.g., “chair”), location constraints (e.g., “bedroom”), and direction constraints (e.g., “turn left”). Particularly, object and location constraints describe the landmarks that can be observed along the desired navigation path. Thus, they will prompt subsequent building of the value map (§ III-C). In practice, we implement the above decomposition and constraints extraction process through an LLM. The prompt consists of four parts, namely task description, output definition, few-shot prompt, and key content reminder. We find that it’s helpful to give large language models an example to follow. Because the few-shot prompt can set a clear expectation of the desired output.
@@ III-C Constraint-aware Value Mapper
Constraint-aware Value Map Generation. The constraint-aware value map generation process first creates a semantic map [71] that captures the environment’s layout and then projects navigation values onto the navigable areas. As shown in Figure 3, we first build a semantic map and extract navigable areas using egocentric RGB-D images and the camera’s pose. Then we use BLIP-2  [5], a pre-trained VLM to compute the cosine similarity value between the current RGB observation and the constraint prompt identified by CSM (§ III-B). The value measures how relevant the current HFOV is for satisfying the constraint prompt. These values are then updated onto the navigable area, forming the value map.
[…]
$c_{\textrm{i,j}}=\textrm{cos}^{2}(\theta/(\theta_{\textrm{hfov}}/2)\cdot\pi/2)$ (1)
where $\theta$ is the angle between the pixel and the optical axis, $\theta_{\textrm{hfov}}$ is a constant angle HFOV. The process is formulated as follows:
$v_{\textrm{i,j}}^{\textrm{curr}}$ $=(c_{\textrm{i,j}}^{\textrm{curr}}v_{\textrm{i,j}}^{\textrm{curr}}+c_{\textrm{i,j}}^{\textrm{prev}}v_{\textrm{i,j}}^{\textrm{prev}})/(c_{\textrm{i,j}}^{\textrm{curr}}+c_{\textrm{i,j}}^{\textrm{prev}});$ (2)
$c_{\textrm{i,j}}^{\textrm{curr}}$ $=\left[(c_{\textrm{i,j}}^{\textrm{curr}})^{2}+(c_{\textrm{i,j}}^{\textrm{prev}})^{2}\right]/(c_{\textrm{i,j}}^{\textrm{curr}}+c_{\textrm{i,j}}^{\textrm{prev}})$
[…]
@@
It is worth noting that when CSM switches constraints, the constraint prompt changes. If the previous value map is cleared entirely, the agent loses navigational cues and must rebuild from scratch, often causing it to linger and collide. To address this, we introduce a historical decay factor $\gamma$ that retains past value map $\mathbf{V}_{t}$ with exponentially reduced weights, helping the agent focus on new constraints while still leveraging past exploration:
@@
$\mathbf{V}_{\textrm{t+1}}=\left\{\begin{aligned} &\gamma\cdot f(\mathbf{V}_{\textrm{t}}),~{}\textrm{if}~{}\textrm{switch constraint}\\ &f(\mathbf{V}_{\textrm{t}})\ \quad,~{}\text{otherwise}\end{aligned}\right.$ (3)
where $f$ denotes the update function for the value map.
@@
To encourage exploration, we introduce a trajectory mask representing the agent’s willingness to explore. This mask starts as a matrix of ones, but its values decay exponentially by a factor of $\lambda$ in regions the agent has already traversed. The value map is then adjusted by element-wise multiplication with this trajectory mask.
@@
Sasra [78] - 8.32 24.0 - 22.0
Seq2Seq [4] - 7.77 25.0 37.0 22.0
AG-CMTP [79] - 7.90 23.1 39.2 19.1
@@
[…]
Sim2Sim [52] - 6.07 43.0 52.0 36.0
GridMM [48] - 5.11 49.0 61.0 41.0
ETPNav [3] - 4.71 57.0 65.0 49.0
BEVBert [53] - 4.57 59.0 67.0 50.0
WS-MGMap [47] - 6.28 38.9 47.6 34.3
NavGPT-CE∗† [7] 8.37 16.3 26.9 10.2
OpenNav (LLama3.1) [9] 7.25 16.0 23.0 12.9
InstructNav (ego) [67] 9.20 17.0 47.0 11.0
A2Nav [66] $\bm{\smallsetminus}$ 22.6 - 11.1
CA-Nav 7.58 25.3 48.0 10.8
IV Experiments And Results
IV-A Experiment Setup
Dataset and Evaluation. We conduct experiments using Habitat simulator [80] on the val-unseen split of R2R-CE [4] and RxR-CE [81], the only two VLN-CE datasets. They provide 1839 and 3669 step-by-step trajectory-instruction pairs across 11 val unseen environments, respectively, with RxR containing more detailed instructions.
We use standard metrics following [1, 4]: Navigation Error (NE), i.e., the mean distance from the final location to the destination, Success Rate (SR), i.e., the proportion of episodes with NE under 3 meters, Oracle Success Rate (OSR), i.e., SR with an oracle stop policy, Success weighted by Trajectory Length (SPL), i.e., SR normalized by trajectory length, Normalized Dynamic Time Warping (NDTW), i.e., the fidelity between the agent’s and the annotated trajectories, and Success-weighted Dynamic Time Warping (SDTW), i.e., NDTW weighted by SR.
IV-B Main Results
[…]
@@ IV-B Main Results
However, CA-Nav exhibits a slightly lower SPL compared to some other methods. This is primarily because, in the egocentric setting, the agent receives more limited observations compared to panoramic views. As a result, the agent spends more steps exploring the environment, which slightly reduces path efficiency. In contrast, OpenNav utilizes panoramic observations along with a waypoint predictor, providing additional information and thereby achieving a higher SPL. Similarly, InstructNav achieves an SPL of 24 with panoramic views, but it drops to 11 when using egocentric observations. Therefore, under the egocentric observation setting, CA-Nav achieves a comparable SPL to other methods.
RxR-CE Dataset. Table II presents the results of the RxR dataset, which contains longer and more detailed instructions with fine-grained descriptions of landmarks and actions. These characteristics of the dataset make it more challenging, as they lead to more frequent instruction switches and more complex constraint identification. Despite these challenges, CA-Nav demonstrates competitive performance, surpassing several models trained with panoramic observations such as LAW or egocentric observations such as WS-MGMap. Notably, CA-Nav outperforms A2Nav in both SR and NE, while remaining competitive in SPL. Although A2Nav pretrained five action-aware navigators it shows lower performance in SR and NE, indicating that CA-Nav’s CSM and CSV are effective even when facing complex and lengthy instructions.
IV-C Ablation Study
We conduct ablation experiments on the R2R dataset to evaluate each component of CA-Nav, focusing on six aspects: (1) the impact of constraints, (2) value map update methods, (3) the effectiveness of superpixel-based waypoint selection, (4) the impact of different superpixel size, (5) robustness to different constraint switching step threshold and (6) generalization to different LLMs.
TABLE III: The effect of different constraints.
@@ IV-C Ablation Study
w/o direction constraint 7.74 24.0 46.9 10.9
w/o object constraint 8.10 20.9 45.4 7.2
w/o location constraint 7.93 23.1 46.6 8.6
w/o all constraints 7.95 20.0 36.4 9.8
w/ all constraints 7.58 25.3 48.0 10.8
@@
How can the value map be better updated? Recall that in the value map generation section § III-C, we designed a historical decay mechanism to leverage past explorations and a trajectory mask to encourage exploration. We investigate their influence in Table IV. In Row1, we ablate both historical decay and trajectory mask which means the value map will be completely reset upon constraint switching. It’s not unexpected that this brings a severe performance drop. In Row 2, we ablate the trajectory mask, the performance drops and we observe that the agent tends to get stuck in narrow spaces such as corridors.
TABLE IV: Influence of value map update methods.
@@
None 7.68 22.3 38.9 10.3
w/ trajectory mask 7.65 24.6 41.3 10.9
w/ historical decay 7.57 24.6 45.6 10.8
trajectory + historical 7.58 25.3 48.0 10.8
This suggests that the trajectory mask contributes positively to the agent’s exploration. In Row 3, where the historical decay is ablated and shows a worse performance. This highlights the inefficiency of discarding all previous knowledge after each constraint switches, as it forces the agent to rediscover previously explored areas. However, Row 4, which incorporates both a trajectory mask and long-term value maintenance using historical decay, effectively balances exploration with exploitation and achieves the best performance.
@@
TABLE V: Different waypoint selection methods.
[…]
FBE-based 8.08 21.9 50.2 10.4
Pixel-based 7.87 22.9 42.9 10.4
Superpixel-based 7.58 25.3 48.0 10.8
@@
While a confidence mask is applied to perform cosine-weighted averaging during value map generation (§ III-C), the trajectory mask affects the map’s smoothness, and the historical decay further disrupts consistency when sub-instructions switch. This indicates the importance of incorporating a more global perspective into the value map utilizing. By clustering similar regions, the superpixel-based method enhances the understanding of the explored area, resulting in more stable navigation planning. This also explains why the superpixel-based method outperforms the Pixel-based approach.
Robustness analysis of superpixel size. The superpixel size determines the granularity of the waypoints used for navigation. Smaller superpixels result in more precise waypoints, while larger superpixels result in coarser regions, which may affect the agent’s ability to make accurate navigation decisions. Therefore, it’s necessary to analyze the impact of different superpixel sizes on CA-Nav’s performance. We test four different superpixel sizes and the resolution ratio was set to 5, which means each pixel corresponds to a 5 cm $\times$ 5 cm area in the real world.
TABLE VI: Different superpixel size.
[…]
$25\times 25$ (pixel) 7.54 23.7 47.5 10.3
$50\times 50$ (pixel) 7.57 24.9 46.5 10.5
$75\times 75$ (pixel) 7.75 24.1 46.0 10.1
$100\times 100$ (pixel) 7.89 21.4 42.7 8.7
@@
The results in Table VI show that a smaller superpixel size (25 $\times$ 25) leads to lower navigation error (NE) and higher object success rate (OSR). This is likely due to the finer granularity of waypoints, which help the agent get closer to the destination. However, the success rate (SR) decreases because smaller superpixels are more sensitive to noise in the value map. On the other hand, when the superpixel size is too large (e.g., 100 $\times$ 100), the regions become overly coarse, making it difficult for CA-Nav to select precise waypoints, resulting in a significant drop in performance. In such cases, the robot may end up in imprecise areas or fail to navigate effectively, particularly when encountering smaller obstacles or narrow paths. Our results suggest that superpixel sizes around 50 $\times$ 50 achieve an optimal balance between precision and robustness to value map noise.
Influence of constraint switching step threshold. Recall that in the sub-instruction switching section§ III-B we introduced minimum and maximum constraint switching step thresholds respectively. This mechanism is designed to ensure that transitions between constraints occur neither too early nor too late, which could negatively affect navigation performance.
TABLE VII: Different constraint switching steps threshold.
[…]
0/0 7.81 22.3 40.2 10.9
5/25 7.73 24.1 45.1 10.9
10/25 7.58 25.3 48.0 10.8
15/25 7.65 24.8 47.6 9.9
10/15 7.54 25.0 48.1 11.7
10/25 7.58 25.3 48.0 10.8
10/35 7.56 24.4 48.5 10.2
@@
By setting appropriate thresholds for the minimum and maximum steps, we can improve the stability of navigation. To figure out its impact on navigation we test various combinations of minimum and maximum thresholds as shown in table VII. As shown in Line 1, When both the minimum and maximum thresholds are set to 0, representing the vanilla constraints transition mechanism, the SR, OSR, and NE all get worse. This demonstrates the importance of the threshold in improving performance. However, CA-Nav shows robustness against minor variations in other thresholds. Whether the minimum is set to 10 or 15, and the maximum to 25 or 35, the performance remains within a narrow range. This suggests that CA-Nav is not overly sensitive to small changes in the constraint switching steps, making it more flexible and adaptable in real-world scenarios.
@@
Generalization to other LLMs. We further replace GPT-4 in our method with GPT-3.5 and Claude 3.5 Sonnet to explore the robustness and generalizability across different LLMs. From Table VIII, we observe that Claude-3.5 Sonnet achieves a comparable SR to GPT-4 and even surpasses it in NE and SPL. This demonstrates that our method can adapt effectively to other LLMs. However, the performance of GPT-3.5 is not satisfactory. The reason could be that GPT-3.5 extracts less precise constraints than GPT-4 and Claude-3.5.
TABLE VIII: Generalization to different LLMs.
[…]
GPT-3.5 7.66 21.1 45.0 9.4
Claude-3.5 Sonnet 7.41 25.2 47.1 11.8
GPT-4 7.58 25.3 48.0 10.8

### v2

@@ II Related Work
Although significant progress has been made, these methods overlook low-level control in continuous environments, particularly in scenarios where predefined navigation graphs are unavailable. As a result, their performance drops significantly when VLN methods are applied to real-world settings [43].
@@ III-B Constraint-aware Sub-instruction Manager
Instruction Decomposition. As shown in Figure 2, at the beginning of an episode, CSM decomposes the instruction into a sequence of sub-instructions. Each of them outlines the goal for the current sub-instruction and specifies the constraints for switching to the next one. We categorize these constraints as object constraints (e.g., “chair”), location constraints (e.g., “bedroom”), and direction constraints (e.g., “turn left”). Particularly, object and location constraints describe the landmarks that can be observed along the desired navigation path. Thus, they will prompt subsequent building of the value map (§ III-C). In practice, we implement the above decomposition and constraints extraction process through an LLM. The prompt consists of four parts, namely task description, output definition, few-shot prompt, and key content reminder (in Appendix Figure 5). We find that it’s helpful to give large language models an example to follow. Because the few-shot prompt can set a clear expectation of the desired output.
@@ III-C Constraint-aware Value Mapper
Constraint-aware Value Map Generation. As navigation progresses, we maintain and continuously update a value map that encodes both the environment layout and instruction-relevant values. Specifically, the environment layout is captured by maintaining a semantic map, from which we derive navigable areas following SemExp [70]. Below, we describe how the instruction-relevant values are projected onto these navigable areas. An overview of this process is illustrated in Figure 4. Recall that in our setting the agent uses egocentric observation space so the visible region is limited by its HFOV. Therefore, the value map $\mathbf{V}_{t}^{\text{global}}$ can be divided into two parts, namely the invisible part $\mathbf{V}_{t}^{\text{invisible}}$ and the visible part $\mathbf{V}_{t}^{\text{visible}}$ where the value update takes place. Formally:
$\mathbf{V}_{t}^{\text{global}}=\left[\left(\mathbf{V}_{t}^{\text{visible}}+\mathbf{V}_{t}^{\text{invisible}}\right)\cdot\gamma^{\mathcal{B}}\right]\odot\mathbf{M}$ (1)
where a decay factor $\gamma$ and a trajectory map $\mathbf{M}$ are include in the global value map update function which will be interpreted in the following parts.
At each time step, given the current RGB observation $\mathbf{O}_{t}$ and the constraint prompt provided by the Constraint Semantic Module (CSM) (§ III-B), we first compute the cosine similarity between them. This similarity score $v_{t}=\text{BLIP2}(\mathbf{O}_{t},\textrm{prompt}_{t})$ reflects how relevant the current observation is to the given instruction constraint. The similarity score $v_{t}$ is applied to the corresponding frustum region and forms the $\mathbf{V_{t}^{\text{visible}}}$. When updating the value map we follow VLFM [71] that utilizes a confidence mask to apply a cosine-weighted average:
$\mathbf{C}_{t}=\left\{\begin{array}[]{ll}\text{cos}^{2}(\theta_{i,j}/(\theta_{\text{hfov}}/2)\cdot\pi/2),&\text{if $(i,j)$ is visible}\\ 0,&\text{if $(i,j)$ is invisible}\end{array}\right.$ (2)
where $\theta_{i,j}$ denotes the angle between pixel $(i,j)$ and the camera’s optical axis, and $\theta_{\text{hfov}}$ is the angle of HFOV. The confidence mask $\mathbf{C}$ assigns higher weights to regions near the optical axis, prioritizing areas with more reliable observations, while assigning zero weight to regions outside the frustum.
We then update the value map following Equation 3, the function $\mathcal{T}$ extracts the frustum region from the previous step of global value map according to current agent pose $\mathbf{p}_{t}$ and forms $\mathbf{V}_{t-1}^{\text{visible}}$, then the confidence mask in visible regions are used to do the cosine-weighted average update.
$\mathbf{V}_{t}^{\text{visible}}$ $\leftarrow\frac{\mathbf{C}_{t}^{\text{visible}}\mathbf{V}_{t}^{\text{visible}}+\mathbf{C}_{t-1}^{\text{visible}}\mathbf{V}_{t-1}^{\text{visible}}}{\mathbf{C}_{t}^{\text{visible}}+\mathbf{C}_{t-1}^{\text{visible}}}$ (3)
$\mathbf{V}_{t-1}^{\text{visible}}$ $\leftarrow\mathcal{T}\left(\mathbf{V}_{t-1}^{\text{global}},\mathbf{p}_{t}\right)$
$\mathbf{V}_{t}^{\text{invisible}}$ $\leftarrow\mathbf{V}_{t-1}^{\text{invisible}};\mathbf{V}_{t}^{\text{invisible}}\text{ is initialized as {0}}$
After updating the value map we update the confidence mask following Equation 4 which is similar to the process of updating value map.
$\mathbf{C}_{t}^{\text{visible}}$ $\leftarrow\frac{\left(\mathbf{C}_{t}^{\text{visible}}\right)^{2}+\left(\mathbf{C}_{t-1}^{\text{visible}}\right)^{2}}{\mathbf{C}_{t}^{\text{visible}}+\mathbf{C}_{t-1}^{\text{visible}}}$ (4)
$\mathbf{C}_{t-1}^{\text{visible}}$ $\leftarrow\mathcal{T}\left(\mathbf{C}_{t-1}^{\text{global}},\mathbf{p}_{t}\right)$
$\mathbf{C}_{t}^{\text{invisible}}$ $\leftarrow\mathbf{C}_{t-1}^{\text{invisible}};\mathbf{C}_{t}^{\text{invisible}}\text{ is initialized as {0}}$
@@
It’s worth noting that when CSM switches constraints, the constraint prompt changes. If the previous value map is cleared entirely, the agent loses navigational cues and must rebuild from scratch, often causing it to linger and collide. To address this, we introduce a historical decay factor $\gamma$ that retains past value map with a reduced weight, helping the agent focus on new constraints while still leveraging past exploration. The historical decay factor is controlled by a binary indicator $\mathcal{B}$ that reflects whether a constraint switch occurs. The decay of the value map happens only when constraint switches. The default $\gamma$ is set to 0.5.
@@
$\gamma\in(0,1];\mathcal{B}=\left\{\begin{array}[]{ll}1,\text{switch constraint}\\ 0,\text{not switch constraint}\end{array}\right.$ (5)
@@
Finally, to encourage exploration, we introduce a trajectory mask $\mathbf{M}$ representing the agent’s willingness to explore, as shown in Equation 6. This mask starts as a matrix of ones, but its values decay exponentially by a factor of $\lambda$ in regions the agent has already traversed, $k(i,j)$ represents the total number of visits to position $(i,j)$. The value map will adjusted by element-wise multiplication with this trajectory mas. By default, the $\lambda$ is set to 0.95.
$\mathbf{M}_{i,j}=\lambda^{k(i,j)},\lambda\in(0,1],\mathbf{M}\text{ is initialized as {1}}$ (6)
@@
IV Experiments And Results
IV-A Experiment Setup
Dataset and Evaluation. We primarily conduct experiments on the val-unseen split of R2R-CE [4] and RxR-CE [77] using Habitat simulator [78], which are the only two VLN-CE datasets. R2R-CE and RxR-CE datasets provide 1839 and 3669 step-by-step trajectory-instruction pairs across 11 unseen environments, respectively, with RxR-CE containing more detailed instructions and longer trajectories.
We use standard metrics following [1, 4]: Navigation Error (NE) measures the mean distance from the final location to the destination; Success Rate (SR) represents the proportion of episodes with NE under 3 meters; Oracle Success Rate (OSR) calculates SR with an oracle stop policy; Success weighted by Trajectory Length (SPL) normalizes SR by trajectory length; Normalized Dynamic Time Warping (NDTW) evaluates the fidelity between the agent’s and the annotated trajectories; and Success-weighted Dynamic Time Warping (SDTW) weights NDTW by SR. For R2R-CE dataset, SR and SPL are the primary metrics to evaluate navigation performance. RxR-CE additionally emphasizes NDTW and SDTW metrics to assess path fidelity.
IV-B Main Results
[…]
NavGPT was originally for zero-shot VLN, we transfer it to zero-shot VLN-CE by using a waypoint predictor [50] to predict navigable nodes in panoramas, discretizing the continuous environment for high-level navigation planning. We also list other methods that rely on high-level actions, including Sim2Sim [52], GridMM [48], ETPNav [3], and BEVBert [53].
@@ IV-B Main Results
Both the detector-based and gradient-based variants of CoW perform poorly, which is unsurprising. The main reason is that CoW’s single-object searching capability is inherently limited [^2] CoW-CLIP-Grad achieved 9.2 SR and CoW-OWL achieved 4.7 SR in Habitat MP3D.. The performance further degrades in VLN-CE, where the task involves multi-object search and is subject to accumulating navigation errors. Moreover, CoW is designed for label-level object navigation (i.e., finding any instance of a given object category), while VLN-CE requires instance-level understanding - the agent needs to locate specific objects or locations that match detailed language descriptions, rather than just finding any object of a certain category.
@@
However, CA-Nav exhibits a slightly lower SPL compared to some other methods. This is primarily because, in the egocentric setting, the agent receives more limited observations compared to panoramic views. As a result, the agent spends more steps exploring the environment, which slightly reduces path efficiency. In contrast, OpenNav utilizes panoramic observations along with a waypoint predictor, providing additional information and thereby achieving a higher SPL. Similarly, InstructNav achieves an SPL of 24 with panoramic views, but it drops to 11 when using egocentric observations. Therefore, under the egocentric observation setting, CA-Nav achieves a comparable SPL to other methods. We also report CA-Nav’s performance on the test-unseen split [^4] To our knowledge, no other zero-shot VLN-CE methods have reported results on the test unseen split., which demonstrates a comparable SR and a slightly higher SPL compared to its performance on the val-unseen split. This indicates that CA-Nav maintains its navigation effectiveness in previously unseen environments. Moreover, the improvement in SPL also highlights its ability to generalize to new instructions and adapt to unfamiliar spatial layouts.
@@
RxR-CE Dataset. Table II reports the results on the RxR-CE val-unseen split [^5] The RxR-CE test-unseen leaderboard has been closed, and the corresponding results are currently unavailable., which contains longer and more detailed instructions with fine-grained descriptions of landmarks and actions. These characteristics of the dataset make it more challenging, as they lead to more frequent instruction switches and more complex constraint identification. Despite these challenges, CA-Nav demonstrates competitive performance, surpassing several models trained with panoramic observations such as LAW or egocentric observations such as WS-MGMap. Notably, CA-Nav outperforms A2Nav in both SR and NE, while remaining competitive in SPL. Although A2Nav pretrained five action-aware navigators it shows lower performance in SR and NE, indicating that CA-Nav’s CSM and CSV are effective even when facing complex instructions.
IV-C Ablation Study
We conduct ablation experiments on the R2R-CE dataset to evaluate each component of CA-Nav, focusing on six aspects: (1) the impact of constraints, (2) value map update methods, (3) the effectiveness of superpixel-based waypoint selection, (4) the impact of different superpixel size, (5) robustness to different constraint switching step thresholds, (6) analysis of different history decay factors, (7) influence of trajectory decay factors and (8) generalization to different LLMs.
@@ IV-C Ablation Study
This further supports our analysis in § I that the agent relies more on landmarks than directions during navigation, which also explains why CA-Nav outperforms A2Nav. Finally, when both object. By designing appropriate constraints and using a Constraint-aware Sub-instruction Manager, our method can autonomously switch between sub-instructions in long-horizon navigation tasks. This not only compensates for VLFM’s [71] inability to handle sequential tasks but also avoids frequent calls to LLMs.
How can the value map be better updated? Recall that in the value map generation section § III-C, we designed a historical decay mechanism to leverage past explorations and a trajectory mask to encourage exploration. We investigate their influence in Table III-3. In Row1, we ablate both historical decay and trajectory mask which means the value map will be completely reset upon constraint switching. It’s not unexpected that this brings a severe performance drop. In Row 2, we ablate the trajectory mask, the performance drops and we observe that the agent tends to get stuck in narrow spaces such as corridors. This suggests that the trajectory mask contributes positively to the agent’s exploration. In Row 3, where the historical decay is ablated and shows a worse performance. This highlights the inefficiency of discarding all previous knowledge after each constraint switches, as it forces the agent to rediscover previously explored areas. However, Row 4, which incorporates both a trajectory mask and long-term value maintenance using historical decay, effectively balances exploration with exploitation and achieves the best performance.
TABLE III: Ablation results for constraints, switching step thresholds, value map update methods, waypoint selection methods, history decay factors, trajectory decay factors, superpixel sizes and LLMs.
1. The effect of different constraints 2. Different constraint switching step thresholds
@@
- - - 7.95 36.4 20.0 9.8 - - - - -
- - 7.79 40.4 22.8 10.1 0/0 7.81 40.2 22.3 10.9
- - 7.71 44.0 23.8 10.7 5/25 7.73 45.1 24.1 10.9
- - 7.85 43.6 23.5 10.4 10/25 7.58 48.0 25.3 10.8
- 8.10 45.4 20.9 7.2 15/25 7.65 47.6 24.8 9.9
- 7.93 46.6 23.1 8.6 10/15 7.54 48.1 25.0 11.7
- 7.74 46.9 24.0 10.9 10/25 7.58 48.0 25.3 10.8
7.58 48.0 25.3 10.8 10/35 7.56 48.5 24.4 10.2
3. Influence of value map update methods 4. Different waypoint selection methods
@@
None 7.68 38.9 22.3 10.3 FBE-based 8.08 50.2 21.9 10.4
trajectory mask 7.65 41.3 24.6 10.9 Pixel-based 7.87 42.9 22.9 10.4
historical decay 7.57 45.6 24.6 10.8 ORP-based 7.54 45.7 25.0 10.6
trajectory mask + historical decay 7.58 48.0 25.3 10.8 Superpixel-based 7.58 48.0 25.3 10.8
5. Analysis of different history decay factors 6. Influence of trajectory decay factors
@@
0.0 7.57 45.5 24.4 10.7 0.95 7.58 48.0 25.3 10.8
0.25 7.93 47.2 24.8 10.7 0.90 7.52 47.9 24.5 10.8
0.50 7.58 48.0 25.3 10.8 0.85 7.57 48.8 24.7 10.8
0.75 7.77 46.0 24.0 10.5 0.80 7.82 46.8 24.8 10.2
7. Different superpixel sizes 8. Generalization to different LLMs
@@
$25\times 25$ (pixel) 7.54 47.5 23.7 10.3 GPT-3.5 7.66 45.0 21.1 9.4
$50\times 50$ (pixel) 7.57 46.5 24.9 10.5 Claude-3.5 Sonnet 7.41 47.1 25.2 11.8
$75\times 75$ (pixel) 7.75 46.0 24.1 10.1 GPT-4 7.58 48.0 25.3 10.8
$100\times 100$ (pixel) 7.89 42.7 21.4 8.7 - - - - -
Comparison of different waypoint selection methods. To verify the effectiveness of the superpixel-based waypoint selection method, we compare it with the FBE-based method mentioned in § III-C. Additionally, we also compare it to the Pixel-based and ORP-based methods. The Pixel-based method directly selects the pixel with the highest value from the global value map as the navigation target. While the ORP-based method chooses the Optimal Region first and then selects the Pixel with the highest value as the waypoint.
@@
Results in Table III-4 show that the superpixel-based method outperforms the FBE-based and Pixel-based methods in terms of SR, and also takes the lead in NE and SPL. This is because value map updates tend to be uneven, making the FBE-based method, which focuses on local frontiers more susceptible to disruptions from sudden value changes (§ 8). While a confidence mask is applied to perform cosine-weighted averaging during value map generation (§ III-C), the trajectory mask affects the map’s smoothness, and the historical decay further disrupts consistency when sub-instructions switch. This indicates the importance of incorporating a more global perspective into the value map utilizing. The ORP-based method performs closely with the superpixel-based method and is superior to FBE-based and Pixel-based methods. This indicates that clustering similar regions would improve the understanding of the explored area, resulting in more stable navigation planning.
@@
Robustness analysis of superpixel size. The superpixel size determines the granularity of the waypoints used for navigation. Smaller superpixels result in more precise waypoints, while larger superpixels result in coarser regions, which may affect the agent’s ability to make accurate navigation decisions. Therefore, it’s necessary to analyze the impact of different superpixel sizes on CA-Nav’s performance. We test four different superpixel sizes and the resolution ratio was set to 5, which means each pixel corresponds to a 5 cm $\times$ 5 cm area in the real world. The results in Table III-7 show that a smaller superpixel size (25 $\times$ 25) leads to lower navigation error (NE) and higher object success rate (OSR). This is likely due to the finer granularity of waypoints, which help the agent get closer to the destination. However, the success rate (SR) decreases because smaller superpixels are more sensitive to noise in the value map. On the other hand, when the superpixel size is too large (e.g., 100 $\times$ 100), the regions become overly coarse, making it difficult for CA-Nav to select precise waypoints, resulting in a significant drop in performance. In such cases, the robot may end up in imprecise areas or fail to navigate effectively, particularly when encountering smaller obstacles or narrow paths. Our results suggest that superpixel sizes around 50 $\times$ 50 achieve an optimal balance between precision and robustness to value map noise. We finally set the default superpixel size as 48 $\times$ 48 which achieves the best result.
@@
Influence of constraint switching step threshold. Recall that in the sub-instruction switching section§ III-B we introduced minimum and maximum constraint switching step thresholds respectively. This mechanism is designed to ensure that transitions between constraints occur neither too early nor too late, which could negatively affect navigation performance. By setting appropriate thresholds for the minimum and maximum steps, we can improve the stability of navigation. To figure out its impact on navigation we test various combinations of minimum and maximum thresholds as shown in table III-2. As shown in Line 1, When both the minimum and maximum thresholds are set to 0, representing the vanilla constraints transition mechanism, the SR, OSR, and NE all get worse. This demonstrates the importance of the threshold in improving performance. However, CA-Nav shows robustness against minor variations in other thresholds. Whether the minimum is set to 10 or 15, and the maximum to 25 or 35, the performance remains within a narrow range. This suggests that CA-Nav is not overly sensitive to small changes in the constraint switching steps, making it more flexible and adaptable in real-world scenarios.
Analysis of different history decay factor. We investigate the impact of the historical decay factor $\gamma$ on final performance in Table III-5. The default setting of $\gamma=0.5$ means that when switching constraints, the historical value map is decayed by half. Both decreasing and increasing $\gamma$ lead to a drop in SR, though the changes are moderate, and the SPL remains stable. Overall, a medium decay factor (e.g., 0.5) proves to be optimal, striking a balance between preserving historical information and adapting to new constraints.
Influence of trajectory decay factor. The influence of trajectory decay factor is also examined, as shown in Table III-6. Specifically, the decay factor $\lambda$ is applied to the trajectory mask to encourage exploration, and it is set to 0.95 by default. We evaluate the performance of CA-Nav under different $\lambda$ values. As $\lambda$ decreases, both SR and OSR exhibit a moderate decline, indicating that CA-Nav remains robust to minor variations in $\lambda$.
@@
Generalization to other LLMs. We further replace GPT-4 in our method with GPT-3.5 and Claude 3.5 Sonnet to explore the robustness and generalizability across different LLMs. From Table III-8, we observe that Claude-3.5 Sonnet achieves a comparable SR to GPT-4 and even surpasses it in NE and SPL. This demonstrates that our method can adapt effectively to other LLMs. However, the performance of GPT-3.5 is not satisfactory. The reason could be that GPT-3.5 extracts less precise constraints than GPT-4 and Claude-3.5.